OpenStudy (anonymous):

Graph the linear equation. 8x − 3y = 0

OpenStudy (anonymous):

(see graph)

OpenStudy (anonymous):

what points did you use?

OpenStudy (anonymous):

i used the origin and its slope, 8/3 RISE 8 RUN 3 y = 8/3x + 0

OpenStudy (radar):

You could also use two points obtained from the 8x-3y=0 By assigning 3 to x. then y =8 so one point is (3,8) assigning 6 to x would result in y=16, 2nd point is 6, 16)